OpenStudy (anonymous):

Which property is demonstrated? (25 + 84) + 16 = 25 + (84 + 16) commutative property of addition commutative property of multiplication associative property of addition associative property of multiplication

OpenStudy (compassionate):

Associative of addition. The order in which it is added does not matter.

OpenStudy (compassionate):

For addition, the rule is "a + (b + c) = (a + b)

OpenStudy (unklerhaukus):

associativity

OpenStudy (unklerhaukus):

\[a + (b + c) = (a + b) + c\]